OpenStudy (anonymous):

solve by substitution method x+6y=50 -5x+5y+15

OpenStudy (anonymous):

your second equation is incorrect

OpenStudy (anonymous):

-5x+5y=-5

OpenStudy (anonymous):

whoops, lost in translation

OpenStudy (anonymous):

x+6y=50 x=50-6y------eqn 1 now, -5(50-6y) +5y=-5 -250+30y+5y=-5 35y=245 y=7 then, x=50-6*7 x=8

OpenStudy (anonymous):

just plug in x = 50 -6y in second eqn - 5(50 -6y) + 5 y = -5 -50 +6y +y = -1 7y = 49 y = 7 so, x = 50 -6y =50 - 42 = 8

OpenStudy (anonymous):

y=7 x=8